OpenStudy (anonymous):

State whether the slope of the line passing through (2, 4) and (5, 2) is positive, negative, zero, or undefined. A. positive B. negative C. zero D. undefined

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Wait wouldn't it be positive since they aren't negative numbers?

OpenStudy (anonymous):

because m = (2-4)/(5-2) = -2/3

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

Slope is about the change, positive and negative numbers have nothing to do with it

OpenStudy (anonymous):

Plus if you graph it out. You can see that the line is descending when you read the graph left to right.